How do i solve this multi-step equation?



4 + x -2(x-6)= 8

Answer by bonster(299)

4 + x -2(x-6)= 8

First look at "-2(x-6)"
Multiply -2 by x and -2 with -6
=-2(x)+(-2)(-6)
=-2x+12

put it back in:
4+x-2(x-6)=8
4+x-2x+12=8
16-x=8

add x to both sides
(16-x)+x=(8)+x
16=8+x

subtract 8 from both sides
(16)-8=(8+x)-8
8=x

x=8